OpenStudy (anonymous):

What is the solution to l x-3 l=7 A. x= -10, 4 B. x=-10, 10 C. x=-4, 4 D. x= -4, 10

OpenStudy (eyust707):



OpenStudy (anonymous):

B; |-10-3| = 7 becomes 10-3 = 7 which is true and |10-3| = 7 becomes 10-3 = 7 which is also true

OpenStudy (eyust707):

Whats in here could either be +7 or -7and it would still give you 7